OpenStudy (anonymous):

Seperate 72 into two parts, one of which shall be 1/3 of the other

OpenStudy (anonymous):

Show all steps

OpenStudy (anonymous):

\[ x+y = 72\] \[y = \frac{1}{3}x\]

OpenStudy (anonymous):

\[\frac{4x}{3} = 72\]

OpenStudy (anonymous):

\[Now \:\:Solve\:\:\]

OpenStudy (anonymous):

The solving part is what I am confused about!!!

OpenStudy (anonymous):

x= 54, 1/3X=18 so 54 and 18

OpenStudy (anonymous):

Thanks